solve 7/9X+7/6=-5/6
Found 2 solutions by jim_thompson5910, mkm123:
Answer by jim_thompson5910(35256)

What do you want to do here? Factor? If so, then...

15a-25b+20 = 5(3a - 5b + 4)

Answer by mkm123(1)

7/9X+7/6=-5/6
sub -7/6 -7/6
____________
7/9X= -2

-9/7 times-7/9X = -2 times -9/7
then 9/7 cancel by 7/9 when its cancel its should be left 2 times 9/7 after that all have to do is 2 times -9 = -18 then divided by 7 the answer should be -18/7